33-9(5-m)=4(m+7) can this answer be a negative decimal?
trying to work out but getting stuck

Answer by checkley71(8403)

33-9(5-M)=4(M+7)
33-45+9M=4M+28
9M-4M=28-33+45
5M=40
M=40/5
M=8
PROOF
33-9(5-8)=4(8+7)
33-9(-3)=4(15)
33+27=60
60=60